OpenStudy (anonymous):

Find the second derivative of x^2y=3

OpenStudy (anonymous):

\[x ^{2}*y=3\]

OpenStudy (zzr0ck3r):

implicite?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

I got 6y/x^2 by the way. No \[\frac{ d ^{2}y }{ dx ^{2} }\]

OpenStudy (zzr0ck3r):

Product rule \(=(x^2)'y+x^2y'=0\\2xy+x^2y'=0\\\) now for the second \((2y+2xy')+2xy'+x^2y''=0\)

OpenStudy (zzr0ck3r):

\(2y+4xy'+x^2y''=0\\y''=\frac{-2y-4xy'}{x^2}=\frac{-2y-4x(\frac{-2y}{x})}{x^2}=\frac{-2y+8y}{x^2}=\frac{6y}{x^2}\)

OpenStudy (anonymous):

Oh Great. Thanks, that's what needed. Cheers.